Calculate the free energy change for translocating a proton out of the mitochondrial matrix, where pHmatrix = 7.6, pHcytosol = 7.2, membrane potential = 175 mV and T = 37°C. {free energy G = 2.303 RT (pHin – pHout) + ZF(membrane potential)} (F = 96,485 J/V·mol).
Answer #1

Free energy change for translocating a proton from mitochondria→ cytosol.

Free energy of the the system is thermodynamically equivalent to capacity of a system to do work

G = 2.303 RT(pH in +pH out) Z F

R= 8.314 J/mol K , pH matrix= pH in =7.6 , pH cytosol = pH out= 7.2, T= 37°C( 37+273= 310K), membrane potential= 0.175 V, F= 96,485 J/V. mol

G= 2.303×8.314×310(7.6-7.2)+0.175×96,485 J

G= 2374.25+ 16884.875 J

G= 19259.12 J